# Seeing What’s Not There: Spurious Correlation in Multimodal LLMs

### v1

@@ Appendix I SpurLens Validation
To estimate HR Gaps, for each class, we take the images in the COCO superclass excluding the target class as the negative images dataset. For this experiment, we remove three additional classes (“sports ball”, “dining table”, and “keyboard” due to significant dataset error). We compute HR Gaps with the same spurious features and present the results in Table LABEL:tab:coco_hr_natural_classwise_results; we use $K=50$ for GPT-4o-mini due to API costs, and $K=100$ for all other models.
@@
2 bicycle water bottle 8.7 water bottle 10.0 water bottle 9.0 water bottle 27.3
3 car road 26.0 road 17.0 road 9.0 road 42.0
4 motorcycle boot 8.7 glove 13.3 glove 14.0 glove 22.3
5 airplane control tower 3.0 control tower 5.0 sky 2.7 control tower 6.0
6 bus curb 18.0 curb 15.3 curb 18.7 curb 25.3
@@
8 truck hill 11.3 mountain 12.0 fence 8.3 road 22.3
9 boat harbor 18.3 buoy 12.7 harbor 20.0 fishing gear 27.0
@@
13 stop sign tree 10.0 tree 14.3 tree 9.0 tree 18.0
14 parking meter pavement 17.3 pavement 12.3 pavement 13.7 pavement 19.3
15 bench grass 18.0 flowerbed 16.0 grass 18.7 tree 25.3
@@
17 cat collar 4.0 collar 5.0 collar 3.7 collar 5.7
18 dog harness 6.7 harness 5.0 leash 5.0 grass 6.7
19 horse bridle 21.7 bridle 21.3 bridle 24.3 bridle 26.0
20 sheep grass 6.7 pasture 15.0 pasture 9.0 grass 14.7
21 cow pasture 8.7 pasture 8.7 pasture 9.7 pasture 7.7
@@
24 zebra rock 1.0 dirt 1.3 grass 3.0 dirt 3.0
25 giraffe savannah 0.7 barbwire 2.0 savannah 2.3 sun 1.7
27 backpack pen 16.7 laptop 16.0 laptop 16.7 pen 24.3
@@
31 handbag couch 17.3 notebook 14.7 umbrella 18.3 umbrella 23.7
32 tie suit 25.0 suit 35.3 suit 16.3 shirt 30.0
33 suitcase luggage tag 49.5 luggage tag 28.0 luggage tag 40.3 luggage tag 59.0
34 frisbee tree 3.7 tree 7.0 tree 6.0 tree 4.7
35 skis pole 7.7 glove 14.3 glove 10.0 glove 14.3
36 snowboard jacket 12.7 jacket 17.7 slope 11.0 jacket 23.3
37 sports ball court 13.8 jersey 10.7 referee 10.3 referee 15.0
@@
39 baseball bat grass 5.3 foul line 4.7 helmet 16.3 helmet 7.7
@@
41 skateboard rail 2.0 shoe 5.0 shoe 4.0 shoe 4.3
@@
50 spoon salt 9.0 condiment 16.0 chopstick 22.0 table 12.0
51 bowl rice 26.7 table 11.0 rice 19.0 soup 31.0
52 banana coconut 4.7 coconut 6.7 coconut 9.0 coconut 12.0
53 apple table 7.7 knife 5.3 table 10.7 knife 9.0
54 sandwich skewer 12.3 chip 14.3 skewer 11.7 skewer 17.7
55 orange tablecloth 9.7 tablecloth 9.3 tablecloth 8.7 fruit knife 9.7
56 broccoli chicken 14.0 chicken 13.0 chicken 11.3 chicken 14.7
57 carrot soup 10.7 soup 10.3 soup 11.7 knife 13.7
58 hot dog backyard 10.3 onion 7.0 backyard 8.7 onion 6.3
59 pizza cutting board 5.2 cutting board 7.3 cutting board 7.0 spice 3.7
@@
62 chair rug 16.3 lamp 14.0 lamp 23.0 rug 17.7
63 couch coffee table 11.3 coffee table 12.0 pillow 23.0 pillow 18.7
64 potted plant wall 14.7 book 12.3 wall 5.7 curtain 21.7
65 bed nightstand 13.0 nightstand 16.0 nightstand 25.3 nightstand 27.3
67 dining table napkin 23.0 napkin 27.0 napkin 21.3 chair 53.0
70 toilet sanitary bin 10.0 trash can 12.3 sanitary bin 16.3 sanitary bin 12.0
72 tv blanket 39.7 coffee table 31.7 dvd player 40.0 dvd player 42.7
@@
76 keyboard usb drive 23.3 usb drive 10.0 usb drive 24.0 usb drive 29.3
77 cell phone sunglass 9.3 wallet 12.0 sunglass 10.0 sunglass 7.7
@@
80 toaster fork 9.0 kitchen counter 10.7 microwave 12.3 kitchen counter 9.3
81 sink sponge 15.7 countertop 13.0 countertop 14.3 towel 19.3
@@
86 vase flower 23.3 flower 30.0 flower 16.3 flower 24.3
87 scissors marker 9.3 string 5.3 ribbon 6.7 fabric 2.0
@@
89 hair drier comb 2.0 makeup remover 6.3 towel 3.3 basket 10.7
90 toothbrush floss 3.1 floss 9.3 floss 10.7 floss 4.3
@@
2 bicycle helmet 3.7 lamp post 10.3 sidewalk 8.0 sidewalk 6.7
3 car tree 16.7 traffic signs 15.3 fence 9.3 fence 20.7
4 motorcycle billboard 9.0 billboard 10.2 billboard 4.7 barrier 2.0
5 airplane baggage cart 1.3 sky 2.3 luggage 1.3 car 0.7
6 bus building 2.7 crosswalk 12.0 crosswalk 7.7 tree 3.3
7 train bridge 2.0 bridge 3.3 bridge 12.3 bridge 4.0
8 truck fence 7.3 tree 17.0 tree 14.0 sign 7.3
9 boat harbor 11.0 harbor 33.5 dock 24.7 dock 14.7
10 traffic light bus stop 6.7 bicycle rack 8.3 bus stop 7.7 crosswalk 6.0
11 fire hydrant car 8.7 car 6.0 storm drain 16.7 storm drain 8.0
13 stop sign utility pole 9.3 building 5.2 utility pole 14.3 vehicle 2.0
14 parking meter car 1.3 bicycle rack 5.7 bicycle rack 18.0 bicycle rack 7.3
15 bench pillow 3.3 bicycle 9.7 coffee cup 6.3 pillow 8.7
16 bird pond 4.0 pond 7.3 feeder 18.7 path 4.0
17 cat shelf 5.7 furniture 3.8 shelf 8.0 door 4.0
18 dog bed 4.3 bed 9.5 bed 10.7 bed 8.0
19 horse bridle 6.3 bridle 5.3 saddle 19.0 trailer 6.0
20 sheep hill 6.0 hill 5.3 flower 6.7 hill 5.3
21 cow hill 3.0 road 7.0 road 15.7 hill 8.0
22 elephant fruit 2.0 bamboo 0.7 path 1.0 fruit 0.0
23 bear log 0.0 cave 0.3 river 2.0 fish 0.0
24 zebra bush 1.0 antelope 2.0 antelope 1.0 stone 0.0
25 giraffe safari vehicle 1.0 flower 1.5 horizon 2.0 safari vehicle 1.3
27 backpack sand 26.7 sand 17.3 trail 18.3 trail 14.0
28 umbrella raincoat 11.0 raincoat 7.5 flower pot 7.0 dog 4.0
31 handbag notebook 41.7 notebook 26.5 sofa 35.7 sofa 11.3
32 tie suit 16.7 suit 12.8 suit 44.7 suit 17.3
33 suitcase terminal gate 7.0 concrete floor 15.2 travel pillow 22.0 baggage cart 11.3
34 frisbee dog 6.3 dog 5.5 pathway 17.7 kite 2.0
35 skis chairlift 9.3 chairlift 9.8 chairlift 60.0 chair lift 8.0
36 snowboard snow gun 11.7 slope 5.7 snow gun 20.7 backpack 2.7
38 kite field 3.3 horizon 3.7 horizon 7.7 horizon 4.7
39 baseball bat home plate 9.0 ballpark 16.2 home plate 6.0 umpire’s gear 5.3
40 baseball glove dirt 13.0 home plate 14.2 bat 20.3 ballpark 8.0
41 skateboard sticker 7.7 ramp 4.3 ramp 13.3 street 4.0
42 surfboard paddle 8.0 wetsuit 8.8 wetsuit 17.7 fishing boat 4.7
@@
44 bottle counter 8.7 cabinet 27.7 box 10.3 cup 28.0
46 wine glass candle 5.7 cork 13.8 coaster 20.3 cork 5.3
47 cup water 8.7 candle 15.3 laptop 15.3 ice 21.3
48 fork plate 7.3 cheese 32.0 cheese 50.0 cheese 21.3
49 knife cutlery holder 18.7 chopping block 30.8 peeler 40.0 napkin 5.3
50 spoon counter 9.3 chopstick 32.5 chopstick 34.3 napkin 6.0
51 bowl soup 19.7 soup 31.7 soup 50.0 ice cream 25.3
52 banana coconut 15.7 paper towel 6.3 coconut 7.3 wall 10.0
53 apple juicer 5.3 picnic blanket 6.3 crate 7.3 crate 2.0
54 sandwich chip 8.0 chip 19.3 chip 19.0 chip 11.3
55 orange textile 6.7 textile 6.7 leaf 18.3 peeler 8.3
56 broccoli lemon 7.3 onion 6.0 carrot 12.3 pasta 4.0
57 carrot soup 15.3 soup 14.2 soup 18.0 bowl 3.3
58 hot dog backyard 5.3 napkin 4.0 grill 1.0 cheese 6.7
59 pizza cheese shaker 4.0 knife 3.3 cheese shaker 2.3 table 1.3
60 donut coffee 5.0 chocolate 4.0 chocolate 10.3 spoon 2.0
61 cake confetti 7.7 candle 9.3 candle 10.7 candle 4.0
62 chair blanket 13.7 magazine 17.3 blanket 24.0 picture 27.3
63 couch blanket 25.7 coffee table 21.3 coffee table 16.3 coffee table 25.3
64 potted plant table 18.0 candle 9.0 table 29.7 table 16.7
65 bed nightstand 16.7 nightstand 12.3 nightstand 4.3 nightstand 15.3
70 toilet diaper pail 3.3 sink 9.8 sanitary bin 2.3 sink 8.7
72 tv console 4.0 couch 21.8 coffee table 9.0 popcorn bowl 5.3
73 laptop mouse 14.3 sticky notes 16.2 notebook 19.7 mouse 10.0
74 mouse computer 8.3 computer 13.0 paper 9.3 blanket 0.0
75 remote magazine 8.7 coaster 37.3 coffee table 20.3 blanket 4.0
76 keyboard mouse 58.0 usb drive 62.2 usb drive 63.3 usb drive 56.0
77 cell phone wallet 16.0 tablet 21.7 wallet 21.7 book 11.3
78 microwave turntable 7.7 oven 23.7 oven 12.0 fruit bowl 14.0
79 oven apron 21.7 microwave 45.3 baking sheet 21.0 microwave 28.7
80 toaster knife 23.7 knife 20.2 microwave 51.3 knife 22.0
81 sink towel 8.3 paper towel 13.2 dish soap 21.0 trash can 22.0
82 refrigerator magnet 10.3 magnet 15.7 magnet 15.0 magnet 25.3
84 book lamp 9.7 pencil 18.2 pencil 10.7 pencil 6.0
85 clock lamp 13.0 couch 6.0 wall 6.0 battery 2.7
86 vase flower 18.7 flower 42.0 flower 29.7 flower 17.3
87 scissors marker 29.3 paper 4.0 pen 27.0 marker 0.7
88 teddy bear doll 32.0 doll 14.2 doll 71.3 book 4.0
89 hair drier brush 5.7 comb 6.0 shampoo 49.3 brush holder 4.0
90 toothbrush toothpaste 11.3 toothpaste 12.2 toothpaste 16.0 mouthwash 0.7
@@
5 airplane control tower 8.7 sky 13.9 control tower 9.7
6 bus crosswalk 3.0 crosswalk 11.8 sidewalk 5.6
7 train tree 9.4 tree 3.8 baggage 6.7
8 truck hill 6.2 mountain 10.1 mountain 8.7
9 boat harbor 59.0 harbor 57.3 harbor 68.3
10 traffic light pedestrian 18.3 street pole 27.9 street pole 25.6
11 fire hydrant sidewalk 12.0 curb 10.7 storm drain 11.3
13 stop sign crosswalk 12.0 building 3.4 traffic light 10.3
14 parking meter pavement 14.5 car 10.8 light fixture 9.0
15 bench sunglass 25.0 laptop 23.6 sunglass 21.7
@@
17 cat blanket 4.0 furniture 2.4 basket 5.7
@@
20 sheep cloud 12.0 cloud 9.6 cloud 17.0
21 cow animal pen 10.3 hill 7.9 haystack 10.3
22 elephant sand 5.3 fruit 0.0 waterhole 7.0
23 bear tree 5.0 river 4.6 sky 3.6
24 zebra antelope 7.7 antelope 5.7 antelope 4.7
@@
27 backpack water bottle 22.7 water bottle 12.6 water bottle 19.0
28 umbrella lamp post 8.7 bag 18.6 lamp post 12.7
31 handbag water bottle 23.3 couch 12.9 earphone 9.7
32 tie suit 22.7 suit 10.4 suit 56.3
@@
34 frisbee tree 25.7 tree 12.8 tree 28.0
35 skis glove 43.0 glove 55.2 glove 41.7
36 snowboard glove 51.0 glove 52.0 slope 24.0
@@
38 kite cloud 20.3 cloud 17.7 cloud 19.0
@@
40 baseball glove bat 17.7 cleat 27.0 bat 41.3
@@
44 bottle box 4.0 cabinet 30.6 box 13.3
46 wine glass candle 14.0 plate 14.8 cutlery 17.7
47 cup window 8.3 sugar 14.6 kettle 20.0
@@
49 knife cutlery holder 10.0 cutlery holder 30.0 pan 20.0
50 spoon pan 15.3 salt 17.0 chopstick 26.3
51 bowl rice 14.3 candy 21.7 dishcloth 26.7
@@
54 sandwich basket 6.3 chip 19.6 counter 4.7
@@
56 broccoli pan 17.7 pasta 20.9 pan 17.0
57 carrot soup 16.3 fruit 0.0 soup 21.3
58 hot dog grill 16.7 backyard 13.3 grill 8.3
59 pizza light fixture 10.7 anchor 10.7 light fixture 16.7
@@
61 cake mood lights 17.3 serviette 16.5 candle 13.7
62 chair magazine 13.7 clock 15.5 magazine 34.0
63 couch blanket 13.8 coffee cup 18.6 pillow 15.9
64 potted plant wall 5.0 candle 10.0 candle 9.0
65 bed book 16.9 book 17.0 teddy bear 16.9
@@
70 toilet plunger 19.3 wall art 18.6 sanitary bin 13.0
72 tv console 8.3 dvd player 41.6 dvd player 42.7
@@
74 mouse block 4.0 computer 10.7 paper 12.3
75 remote wall 12.0 wall 28.0 speaker 25.3
76 keyboard usb drive 9.7 chair 21.7 sticky notes 11.3
@@
80 toaster window 8.7 wall cabinets 9.7 microwave 28.0
81 sink countertop 26.3 towel 30.4 sponge 26.7
82 refrigerator fruit 8.5 juice 15.7 butter 7.3
84 book pencil 7.7 clock 17.1 pencil 8.3
85 clock book 13.7 table 3.1 chair 2.3
86 vase art pieces 11.3 flower 51.6 flower 40.7
87 scissors marker 16.7 ribbon 6.4 pen 42.7
88 teddy bear toy box 20.3 doll 16.8 doll 47.3
@@
90 toothbrush toothpaste 13.7 toothpaste 22.6 toothpaste 44.7
Appendix N Class-wise SpurLens Results on Imagenet Subset
As described in Sec. 4, the 100 Imagenet classes considered in Spurious Imagenet [20], we use GPT-4 to generate 32 potential spurious features.
@@
0 tench net 10.0 net 16.0 cloud 10.0 gravel 10.0
2 great white shark boat 7.3 fish 18.7 fish 20.7 wave 18.0
14 indigo bunting grass 0.0 branch 6.7 flower 1.3 fruit 7.3
@@
50 American alligator sand 2.0 water 15.3 marsh 34.7 water 17.3
58 water snake rock 3.3 reed 6.7 reed 11.3 reed 6.0

### v2

@@ Appendix I Hyperparameter Sensitivity Analysis
To estimate HR Gaps, for each class, we take the images in the COCO superclass excluding the target class as the negative images dataset. For this experiment, we remove three additional classes (“sports ball", “dining table", and “keyboard") due to significant dataset error. We compute HR Gaps with the same spurious features and present the results in Table LABEL:tab:coco_hr_natural_classwise_results; we use $K=50$ for GPT-4o-mini due to API costs, and $K=100$ for all other models.
@@
2 bicycle basket 16.5 basket 13.7 basket 10.0 water bottle 27.3
3 car road 26.0 road 17.0 pavement markings 9.3 road 42.0
4 motorcycle mirror 12.2 glove 13.3 glove 14.0 glove 22.3
5 airplane control tower 3.0 turbine 7.9 turbine 5.0 turbine 8.4
6 bus curb 18.0 signage 15.7 curb 18.7 curb 25.3
@@
8 truck hitch 43.3 roof rack 23.5 hitch 33.9 hitch 52.4
9 boat oar 19.5 buoy 12.7 dock 18.7 fishing gear 27.0
@@
13 stop sign tree 10.0 tree 14.3 landscape 12.7 tree 18.0
14 parking meter ticket 31.0 ticket 23.7 ticket 20.3 ticket 28.7
15 bench bush 18.7 flowerbed 16.0 book 19.4 tree 25.3
@@
17 cat collar 4.0 collar 5.0 scratching post 4.1 collar 5.7
18 dog collar 21.8 collar 30.2 collar 14.3 collar 28.1
19 horse saddle 9.7 saddle 7.7 saddle 10.0 saddle 10.3
20 sheep path 3.0 pasture 15.0 pasture 9.0 pasture 12.0
21 cow farmer 9.1 barn 13.2 barn 13.3 barn 11.5
@@
24 zebra trail 1.0 dirt 1.3 grass 3.0 dirt 3.0
25 giraffe fence 0.7 barbwire 2.0 savannah 2.3 savannah 1.3
27 backpack pen 16.7 laptop 16.0 key 19.7 pen 24.3
@@
31 handbag bag stand 20.0 lipstick 17.2 key 21.0 wallet 39.8
32 tie lapel pin 33.3 suit 35.3 lapel pin 17.7 pocket square 30.7
33 suitcase baggage cart 36.5 baggage cart 15.7 travel brochure 17.3 baggage cart 28.3
34 frisbee park 4.0 tree 7.0 park 8.3 tree 4.7
35 skis pole 7.7 glove 14.3 snowboard 13.4 glove 14.3
36 snowboard pant 19.2 pant 18.4 terrain park 23.0 pant 23.8
37 sports ball turf 11.3 jersey 10.7 referee 10.3 referee 15.0
@@
39 baseball bat uniform 5.5 uniform 8.8 uniform 27.1 uniform 8.2
@@
41 skateboard wall 2.0 shoe 5.0 shoe 4.0 shoe 4.3
@@
50 spoon bowl 28.5 food 35.4 sugar 25.3 sugar 20.8
51 bowl granola 27.3 granola 15.3 granola 19.7 fruit 34.7
52 banana coconut 4.7 fruit label 9.6 coconut 9.0 coconut 12.0
53 apple table 7.7 knife 5.3 table 10.7 tree 15.3
54 sandwich mayonnaise 18.9 pickle 21.0 pickle 18.8 mayonnaise 23.8
55 orange fruit basket 15.3 fruit basket 10.7 fruit basket 14.7 fruit basket 14.5
56 broccoli chicken 14.0 chicken 13.0 chicken 11.3 quinoa 18.8
57 carrot vegetable 30.5 vegetable 29.0 vegetable 33.6 vegetable 31.5
58 hot dog ketchup 12.2 ketchup 17.7 ketchup 24.4 ketchup 23.7
59 pizza olive 8.8 onion 10.1 olive 14.0 mushroom 12.0
@@
62 chair cushion 26.3 lamp 14.0 table 29.3 table 32.3
63 couch throw pillows 14.7 throw pillows 14.5 throw pillows 36.4 throw pillows 20.1
64 potted plant tag 14.8 floor 15.7 vase 12.6 curtain 21.7
65 bed sheet 23.1 sheet 21.9 sheet 29.4 sheet 34.7
67 dining table napkin 23.0 plate 34.7 centerpiece 21.7 chair 53.0
70 toilet sanitary bin 10.0 shower curtain 9.3 sanitary bin 16.3 sanitary bin 12.0
72 tv dvd player 39.7 coffee table 31.7 sound system 40.0 dvd player 42.7
@@
76 keyboard wrist rest 42.2 wrist rest 19.1 wrist rest 45.7 wrist rest 67.2
77 cell phone sunglass 9.3 watch 9.7 sunglass 10.0 sunglass 7.7
@@
80 toaster cutting board 9.7 kitchen counter 10.7 microwave 12.3 kitchen counter 9.3
81 sink sponge holder 18.4 towel 11.3 sponge holder 13.4 sponge holder 23.6
@@
86 vase leaf 27.3 flower 30.0 leaf 19.7 leaf 33.0
87 scissors marker 9.3 felt 20.6 glue 11.2 felt 10.3
@@
89 hair drier conditioner 6.0 conditioner 12.0 conditioner 5.0 basket 10.7
90 toothbrush comb 10.1 floss 9.3 comb 12.0 comb 6.0
@@
2 bicycle basket 24.0 lamp post 11.3 basket 24.3 sidewalk 8.0
3 car tree 17.3 traffic signs 17.7 tree 17.3 tree 17.3
4 motorcycle mirror 10.0 billboard 9.8 billboard 6.7 barrier 1.3
5 airplane baggage cart 1.3 sky 2.7 luggage 1.3 car 0.7
6 bus traffic light 1.7 traffic cone 12.7 post box 6.3 tree 3.3
7 train graffiti 2.3 bridge 3.3 bridge 12.3 bridge 4.0
8 truck hitch 8.3 roof rack 22.5 parking lot 24.0 sign 9.3
9 boat sail 15.0 dock 28.8 dock 25.0 dock 18.7
10 traffic light bus stop 7.7 crosswalk 10.7 bus stop 7.7 crosswalk 8.7
11 fire hydrant car 9.0 street 10.2 storm drain 18.7 storm drain 9.3
13 stop sign crosswalk 9.7 side street 7.0 utility pole 15.3 side street 4.7
14 parking meter car 2.3 bus stop 8.7 bus stop 18.0 curb 5.3
15 bench trashcan 5.0 umbrella 12.7 coffee cup 6.0 coffee cup 8.0
16 bird pond 3.0 sky 7.8 feeder 18.3 bush 4.0
17 cat tv 5.3 furniture 5.2 shelf 8.3 door 5.3
18 dog collar 3.7 bed 7.5 toy 8.0 bed 8.7
19 horse saddle 6.0 barn 4.0 saddle 19.3 trailer 6.0
20 sheep hill 6.0 hill 4.7 shepherd 6.3 hill 5.3
21 cow farmer 6.7 road 7.0 road 17.3 hill 7.3
22 elephant sky 0.7 bamboo 0.7 path 1.0 mud 0.0
23 bear rock 0.0 cave 0.0 river 1.3 log 0.0
24 zebra bush 1.0 antelope 2.0 antelope 1.0 bird 0.0
25 giraffe safari vehicle 1.0 cloud 0.5 horizon 2.0 safari vehicle 1.3
27 backpack charger 23.0 laptop 20.3 first aid kit 26.7 trail 16.0
28 umbrella raincoat 9.3 raincoat 6.2 flower pot 6.0 dog 2.0
31 handbag notebook 43.3 notebook 25.2 notebook 35.0 sofa 10.7
32 tie suit 14.7 suit 10.0 suit 40.3 suit 15.3
33 suitcase baggage cart 7.3 concrete floor 13.8 travel pillow 22.7 baggage cart 6.7
34 frisbee dog 5.3 dog 5.8 dog 18.7 people 2.0
35 skis snow shovel 10.7 snow shovel 13.3 snow pants 65.0 snow shovel 12.0
36 snowboard goggle 6.7 snow 4.2 goggles 17.7 terrain park 4.0
38 kite field 3.3 horizon 2.7 horizon 7.0 horizon 4.7
39 baseball bat home plate 9.0 ballpark 14.8 home plate 6.0 ballpark 6.0
40 baseball glove uniform 15.3 uniform 16.0 bat 21.0 uniform 9.3
41 skateboard sticker 7.7 ramp 4.0 ramp 13.3 street 4.0
42 surfboard paddle 8.0 wetsuit 9.2 wetsuit 16.7 fishing boat 4.7
@@
44 bottle counter 6.0 wall 28.7 box 8.3 cup 22.7
46 wine glass candle 6.3 cork 12.5 coaster 19.0 cork 5.3
47 cup tea 35.3 tea 33.7 tea 25.0 tea 22.7
48 fork plate 8.0 food 28.5 food 45.3 cheese 20.0
49 knife cutlery holder 21.3 utensil 32.0 utensil 40.0 utensil 12.0
50 spoon whisk 20.0 chopstick 30.2 whisk 36.0 whisk 21.3
51 bowl pudding 17.3 pudding 27.7 vegetable 26.3 ice cream 26.0
52 banana coconut 17.3 paper towel 5.7 coconut 6.3 wall 10.0
53 apple basket of fruit 6.7 picnic blanket 6.7 basket of fruit 12.3 table 2.0
54 sandwich mustard 31.3 mustard 67.7 mustard 66.7 mustard 40.7
55 orange wall 11.3 wall 4.7 wall 20.7 cutting board 9.0
56 broccoli lemon 7.3 lemon 8.0 carrot 11.7 napkin 4.0
57 carrot vegetable 15.7 soup 11.2 vegetable 37.0 bowl 3.3
58 hot dog soda 2.3 cheese 2.8 relish 2.0 cheese 6.7
59 pizza basil 2.7 olive 4.3 knife 2.0 basil 2.0
60 donut coffee 6.0 coffee 2.2 coffee 5.7 table 2.0
61 cake chocolate 9.0 chocolate 12.7 candle 12.0 candle 4.0
62 chair cushion 35.7 footrest 44.7 footrest 52.3 cushion 46.7
63 couch blanket 27.3 television 17.3 coffee table 14.3 coffee table 21.3
64 potted plant vase 51.7 vase 24.3 vase 83.0 vase 36.7
65 bed sheet 19.0 sheet 21.7 sheet 13.0 nightstand 14.7
70 toilet diaper pail 3.3 sink 9.8 sink 2.3 sink 8.7
72 tv game console 4.7 couch 27.7 coffee table 11.7 rug 4.0
73 laptop mouse 13.3 webcam 13.3 chair 13.0 mouse 10.0
74 mouse computer 9.3 computer 13.0 paper 10.0 lab equipment 0.7
75 remote floor 9.3 coaster 28.7 window 22.7 blanket 4.0
77 cell phone charger 7.7 key 10.0 bag 11.7 book 8.7
78 microwave cooking tray 4.0 oven 22.7 oven 9.3 plastic wrap 6.7
79 oven pan 19.7 microwave 44.5 pan 21.7 microwave 28.7
80 toaster knife 23.7 coffee maker 19.7 microwave 50.3 knife 20.0
81 sink trash can 8.7 trash can 11.0 knife 23.0 trash can 18.0
82 refrigerator magnet 13.7 magnet 14.3 magnet 14.3 magnet 25.3
84 book lamp 9.3 note 24.3 pencil 12.7 pencil 4.7
85 clock lamp 10.0 alarm button 9.0 alarm button 7.3 alarm button 6.7
86 vase flower 18.0 flower 38.8 plant 30.3 flower 17.3
87 scissors marker 28.3 cardboard 6.7 glue 35.0 glue 2.0
88 teddy bear cuddly blanket 13.0 bed 6.5 cuddly blanket 14.3 book 4.0
89 hair drier conditioner 10.0 conditioner 5.8 shampoo 40.7 makeup remover 4.0
90 toothbrush toothpaste 11.3 comb 13.8 comb 24.0 shaving cream 4.0
@@
5 airplane turbine 19.6 sky 13.9 turbine 22.3
6 bus parking meter 5.9 crosswalk 11.8 signage 7.0
7 train graffiti 10.1 graffiti 11.1 tunnel 8.0
8 truck hitch 12.6 roof rack 16.2 parking lot 36.5
9 boat fishing gear 26.7 buoy 21.2 buoy 34.4
10 traffic light pedestrian 18.3 crosswalk 21.1 curb 23.3
11 fire hydrant sidewalk 12.0 street 12.2 storm drain 11.3
13 stop sign side street 18.6 side street 7.8 traffic light 10.3
14 parking meter pavement 14.5 curb 10.9 bus stop 20.7
15 bench sunglass 25.0 laptop 23.6 book 28.1
@@
17 cat scratching post 14.3 scratching post 5.7 scratching post 14.0
@@
20 sheep barn 11.3 barn 9.4 hill 10.7
21 cow animal pen 10.3 pasture gate 12.1 bale 13.9
22 elephant sand 5.3 bush 0.0 waterhole 7.0
23 bear tree 5.0 river 4.6 tree 2.3
24 zebra waterhole 12.0 waterhole 16.8 waterhole 5.3
@@
27 backpack first aid kit 26.6 water bottle 12.6 first aid kit 24.0
28 umbrella lamp post 8.7 bag 18.6 people 27.1
31 handbag bag stand 32.9 key 16.2 bag stand 27.5
32 tie suit 22.7 pocket square 14.0 suit 56.3
@@
34 frisbee people 28.2 bicycle 14.0 tree 28.0
35 skis glove 43.0 glove 55.2 snow pants 42.6
36 snowboard glove 51.0 glove 52.0 terrain park 31.7
@@
38 kite backpack 18.0 mountain 16.7 camera 7.3
@@
40 baseball glove bat 17.7 uniform 29.5 bat 41.3
@@
44 bottle wall 9.9 cabinet 30.6 box 13.3
46 wine glass candle 14.0 ice bucket 16.5 ice bucket 18.0
47 cup window 8.3 tea 23.8 tea 23.7
@@
49 knife chef hat 13.4 cutlery holder 30.0 pan 20.0
50 spoon whisk 21.3 food 26.8 chopstick 26.3
51 bowl rice 14.3 yogurt 22.8 dishcloth 26.7
@@
54 sandwich basket 6.3 chip 19.6 cheese 5.9
@@
56 broccoli pan 17.7 pasta 20.9 quinoa 18.4
57 carrot vegetable 26.5 fruit 0.0 vegetable 55.5
58 hot dog grill 16.7 soda 10.2 grill 8.3
59 pizza parchment paper 22.2 light fixture 10.6 light fixture 16.7
@@
61 cake banner 22.3 balloon 17.6 banner 18.3
62 chair footrest 20.7 footrest 16.3 footrest 38.6
63 couch pillow 13.8 coffee cup 18.6 throw pillows 18.1
64 potted plant tag 16.9 tag 20.1 pebble 21.0
65 bed throw blanket 18.4 book 17.0 teddy bear 16.9
@@
70 toilet sanitary bin 13.7 wall art 18.6 sanitary bin 13.0
72 tv sound system 16.6 dvd player 41.6 dvd player 42.7
@@
74 mouse lab equipment 5.3 computer 10.7 lab equipment 16.3
75 remote charger 32.3 wall 28.0 speaker 25.3
76 keyboard wrist rest 11.9 chair 21.7 sticky notes 11.3
@@
80 toaster cutting board 8.0 wall cabinets 9.7 microwave 28.0
81 sink sponge holder 24.4 towel 30.4 sponge 26.7
82 refrigerator milk 11.9 milk 17.7 yogurt 7.8
84 book pencil 7.7 clock 17.1 note 9.2
85 clock lamp 13.7 table 3.1 chair 2.3
86 vase decorative balls 15.6 flower 51.6 flower 40.7
87 scissors marker 16.7 craft box 8.7 craft box 47.1
88 teddy bear toy box 20.3 toy box 16.3 toy box 32.0
@@
90 toothbrush comb 14.3 toothpaste 22.6 toothpaste 44.7
S.3 Imagenet Subset
As described in Section 4, the 100 Imagenet classes considered in Spurious Imagenet (spnet), we use GPT-4 to generate 32 potential spurious features.
@@
0 tench net 10.0 reed 22.0 fishing rod 18.0 reed 5.3
2 great white shark boat 7.3 fish 18.7 fish 20.7 fish 14.0
14 indigo bunting branch 0.0 branch 6.7 stone 2.0 branch 4.7
@@
50 American alligator turtle 4.0 lily pads 24.7 water 21.3 water 17.3
58 water snake cattail 3.3 dragonfly 6.7 pond 30.7 reed 6.0